Gaseous methane CH4  will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 6.3 g of methane is mixed with 15.0 g of oxygen. Calculate the maximum mass of water that could be produced by the chemical reaction. Round your answer to 3 significant digits.




1
Expert's answer
2021-10-08T02:09:09-0400

CH4 + 2O2 → CO2 + 2H2O

M(CH4) = 16.04 g/mol

n(CH4) "= \frac{6.3}{16.04}=0.393 \;mol"

M(O2) = 15.99 g/mol

n(O2) "= \frac{15.0}{15.99}=0.938 \;mol"

For each mole of CH4 we need only 2 moles of O2, but for 0.393 mol of CH4 we have more O2 than needed. So, CH4 is the limiting reactant.

According to the reaction:

n(H2O) = 2n(CH4) "= 2 \times 0.393 = 0.786 \;mol"

M(H2O) = 18.01 g/mol

m(H2O) "= 18.01 \times 0.786 = 14.15 \;g"

Answer: 14.1 g.